OpenStudy (anonymous):

.25(8z - 4) = z + 8 - 2z

OpenStudy (anonymous):

first distribute the .25

OpenStudy (anonymous):

2z - .5 = -1z +8 ?

OpenStudy (anonymous):

close, what is 4*.25 or 4(1/4)?

OpenStudy (anonymous):

everything else in that step was correct

OpenStudy (anonymous):

oHHH

OpenStudy (anonymous):

so what do you get now?

OpenStudy (anonymous):

2z - 1 = -1z + 8

OpenStudy (anonymous):

correct, so the next step is to get the z's and the #'s on the same side

OpenStudy (anonymous):

i subtracted 2z from both sides & got; 1 = -3z +8

OpenStudy (anonymous):

then subtracted 8 from the answer & 8

OpenStudy (anonymous):

i got -7 = -3z ?

OpenStudy (anonymous):

close, there is one small thing wrong with your equation 1=-3z+8 if you fix this, and redo the steps in the same way, you will get a nice whole number for your answer. what do you think your mistake was?

OpenStudy (anonymous):

To make it a little easier, I will put the two equations you typed next to each other. 2z-1=-z+8 1=-3z+8 <--Remember there is something wrong with this equation

OpenStudy (anonymous):

i dont know, i cant figure it out.

OpenStudy (anonymous):

ok, well you have the first equation: 2z-1=-z+8 ~~~you subtracted 2z from each side -2z -2z ~~~ which gives you -1=-3z+8 <-- this is correct. you simply forgot to leave the 1 as negative. So using the correct equation: -1=-3z+8 what do you get as your answer now?

OpenStudy (anonymous):

z = -3

OpenStudy (anonymous):

that is soooooooo close!!! but if you have -3z=-9 what do you divide -9 by to get positive z all by itself?

OpenStudy (anonymous):

wait i messed up on writing my reply, i got 3.

OpenStudy (anonymous):

haha ok, but that is the correct answer!!! Good Job!!

OpenStudy (anonymous):

thank you! could i ask for help on another problem really fast?

OpenStudy (anonymous):

definitely!

OpenStudy (anonymous):

the problem is -12h + 39 = -4h - 17 i know how to get the answer but do i add 4h to -4h to cancel it out or do i still subtract?

OpenStudy (anonymous):

you would add 4h to both sides

OpenStudy (anonymous):

okay i got h=7

OpenStudy (anonymous):

That is correct!

OpenStudy (anonymous):

okay, thank you!